Question (3points) a See page 272 4th attempt See Hint You have been asked by the king of the small island nation of Littleston to present the labor force data listed below Number of employed Working-age population Labor force 20,000 40,000 36000 Using this data, calculate the number of unemployed people, the unemployment rate, and the labor force participation rate Enter your answers in the table below. (NOTE: for the unemployment rate and the labor force participation rate, round to one place past the decimal point.) Number of unemployed people Labor force participation rate 3rd attempt 09:13 > VIEW SOLUTION OF1 DELL
Answer #1

Number of Employed = 20000

Working-age population = 40000

Labor Force = 36000

Number of unemployed = Labor Force - Employed = 36000 - 20000 = 16000

Unemployment Rate = Unemplyment number / Labor Force = 16000/36000 = 0.4444 or 44.44 %

Labor force participation rate = Labor Force / Working Age Population = 36000 / 40000 = 0.9 or 90%